OpenStudy (anonymous):

Integrate: [x/sqrt(1+x)]dx

myininaya (myininaya):

did you try letting u=ugly stuff

myininaya (myininaya):

the ugly stuff being sqrt(1+x)

myininaya (myininaya):

\[u=\sqrt{1+x} => u^2=1+x => 2 u du=dx => 2 du=\frac{dx}{u}=\frac{dx} {\sqrt{1+x}}\] So we have \[\int\limits_{}^{} ( u^2-1) 2du\]

myininaya (myininaya):

I replaced dx/sqrt(1+x) with 2 du

myininaya (myininaya):

and i replace x with u^2-1 since we had u^2=1+x

myininaya (myininaya):

\[2(\frac{u^3}{3}-u)+C\] now just replace u with sqrt(1+x) and you will be done :)

OpenStudy (anonymous):

thanks very much:)